OpenStudy (anonymous):

Which system has (1, 3) as the solution? Which system has infinite solutions? Which system has (-8, 0) as the solution? Which system has no solutions? Which system has (-2, -2) as the solution? A) x + y = -1 4x + 4y = -4 B) x - 4y = -8 2x - 3y = -16 C) x + y = -4 -3x + 2y = 2 D) 4x + 6y = -12 2x + 3y = 9 E) x - y = -2 3x - y = 0

OpenStudy (anonymous):

A) 4x+4y=-4(multi eq 1 with 4) 4x+4y=-4 subtract both eq we'll get 0 = 0

OpenStudy (anonymous):

thats a right ?

OpenStudy (anonymous):

wait thats the 4th one so A= 4

OpenStudy (anonymous):

or a = 2

OpenStudy (anonymous):

B) 2x - 8y = -16(multi eq1 with 2) 2x - 3y = -16 subtracting both eqs -5y = 0 y=0 x= -8 (-8,0)

OpenStudy (anonymous):

A system of eq has no solution

OpenStudy (anonymous):

so A= is the 4th one like i said

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what bout the other two

OpenStudy (anonymous):

can you solve the rest ???

OpenStudy (anonymous):

Which system has (1, 3) as the solution?

OpenStudy (anonymous):

B system of eq has solution set = (-8,0)

OpenStudy (anonymous):

E) x - y = -2 3x - y = 0

OpenStudy (anonymous):

that is correct right ?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

E) x - y = -2 3x - y = 0 subtracting both eqs -2x=-2 x= 1 y=3 (1,3) yes E has solution set (1,3)

OpenStudy (anonymous):

Thought so..

OpenStudy (anonymous):

is c the last option

OpenStudy (anonymous):

C)2x +2 y = -8(multi with 2) -3x + 2y = 2 subtract both eqs 5x=-10 x= -2 y=-2 (-2,-2) C has solution set (-2,-2)

OpenStudy (anonymous):

so i am right :)

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

then D is the second one

OpenStudy (anonymous):

D) 4x + 6y = -12 4x + 6y = 18(multi by 2) subtract both eqs 0 = -30 so D has infinite solutions